OpenStudy (anonymous):

Jen made several recipes and used a total of 8\frac{3}{4} cup of flour. If she only had a \frac{1}{4} cup measuring cup, how many times did she have to fill it. A. 35 B. 32 C. 8 D. 11

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

C obviously?

OpenStudy (anonymous):

C or D

OpenStudy (alexandervonhumboldt2):

\[8\frac{ 3 }{ 4 } / \frac{ 1 }{ 4 }=\frac{ 35}{ 4 }*4=35\]

OpenStudy (anonymous):

XD

OpenStudy (alexandervonhumboldt2):

everything is correct lol

OpenStudy (anonymous):

35 times? What the heck

OpenStudy (alexandervonhumboldt2):

1/4*35=35/4=8 3/4

OpenStudy (alexandervonhumboldt2):

35 is the truth

OpenStudy (anonymous):

XD